OpenStudy (anonymous):

The Sum of Three numbers is 123. the second number is 9 less than 2 times the 1st number. the 3rd number is 6 more than 3 times the 1st number. find the 3 numbers.

OpenStudy (anonymous):

Let the 1st number be x Let the 2nd number be y Let the 3rd number be z y=(2x)-9 z=(3x)+6 We know x+y+z=123 Now we add up the 3 numbers [x+(2x-9)+3x+6)=123 6x-3=123 6x=126 x=21 y=33 z=69

OpenStudy (gw2011):

Let x=!st number 2x-9=2nd number 3x+6=3rd number x+(2x-9)+(3x+6)=123 x+2x-9+3x+6=123 6x-3=123 6x=126 x=21 This is the 1st number 2x-9=2(21)-9=33 This is the 2nd number 3x+6=3(21)+6=69 This is the 3rd number